OpenStudy (anonymous):

Can someone help me simplify this- step by step -5+i/2i

OpenStudy (misty1212):

HI!!

OpenStudy (trojanpoem):

Multiply times the conjunct of the denominator: the conjunct of an imaginary number a + bi is a - bi so we have -2i as the conjunct 10 i - 2i^2 / -4i^2 i^2 = 1 10i - 2

OpenStudy (anonymous):

So 10i-2 is correct? I was kind of thinking it was 1/2+5i/2

OpenStudy (anonymous):

@TrojanPoem ^

OpenStudy (trojanpoem):

Sorry, You're totally correct. i = sqrt(-1) i^2 = -1 -5 + i /2i -2i ( -5 + i) / -4i^2 10i - 2i^2 / -4i^2 10i + 2/ 4 1/2 + 5i/2 I repeat my excuse, maybe I rushed a bit.